OpenStudy (anonymous):

Simplify 2 * x/(x+3) * (x-7)/(3x) thanks! :)

OpenStudy (campbell_st):

multiply numerators and denominators which will result in \[\frac{2 \times x \times (x -7)}{(x + 3)\times 3x}\] remove the common factor x, whats remaining is the answer